# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ Abstract.
Let $U$ be a set of positive roots of type $ADE$, and let $\Omega_{U}$ be the set of all maximum cardinality orthogonal subsets of $U$. For each element $R\in\Omega_{U}$, we define a generalized Rothe diagram whose cardinality we call the level, $\rho(R)$, of $R$. We define the generalized quantum Hafnian of $U$ to be the generating function of $\rho$, regarded as a $q$-polynomial in $U$. Several widely studied algebraic and combinatorial objects arise as special cases of these constructions, and in many cases, $\Omega_{U}$ has the structure of a graded partially ordered set with rank function $\rho$. A motivating example of the construction involves a certain set of $k^{2}$ roots in type $D_{2k}$, where the elements of $\Omega_{U}$ correspond to permutations in $S_{k}$, the generalized Rothe diagrams are the traditional Rothe diagrams associated to permutations, the level of a permutation is its length, the generalized quantum Hafnian is the $q$-permanent, and the partial order is the Bruhat order. We exhibit many other natural examples of this construction, including one involving perfect matchings, two involving labelled Fano planes, and one involving the invariant cubic form in type $E_{6}$.
@@
root system, Weyl group, Rothe diagram, quantum Hafnian, quasiparabolic set
@@ 1. Introduction
Let $U$ be a subset of the positive roots of a Weyl group of type $ADE$, and let $\Omega_{U}$ be the set of orthogonal subsets of $U$ of maximal cardinality. For each $R\in\Omega_{U}$, we define a residue of $R$ with respect to $U$ to be an element $\gamma\in U$ such that $s_{\beta}(\gamma)$ is positive for all $\beta\in R$, where $s_{\beta}$ denotes the reflection associated to the vector $\beta$. The generalized Rothe diagram of $R$ is the set of residues of $R$ in $U$, which we denote by $\text{\rm Res}_{U}(R)$. We call $\rho(R):=|\text{\rm Res}_{U}(R)|$ the level of $R$, and we define the generalized quantum Hafnian of $U$ to be the element of ${\mathbb{Z}}[U][q]$ given by
@@
Although it is not a priori clear why one might want to study these definitions, which are new to the best of our knowledge, we will show that many choices for $U$ give rise to concise constructions of known algebraic and combinatorial objects, in some cases endowing them with additional structure. For example, we will use our methods to associate a generalized Rothe diagram in a canonical way to each of the $45$ terms of the invariant cubic polynomial in type $E_{6}$, which gives a nonrecursive combinatorial interpretation of the coefficients of the polynomial.
One of the main motivating examples of this phenomenon (Theorem 3.13) is the case where $U=\{\varepsilon_{i}+\varepsilon_{k+j}:1\leq i,j\leq k\}$ is a set of $k^{2}$ roots of type $D_{2k}$, corresponding to the positions $(i,j)$ on a $k\times k$ grid. Here, the set $\Omega_{U}$ corresponds to the set of configurations of $k$ nonattacking rooks, or equivalently, the elements of the symmetric group $S_{k}$. If we place the origin at the top left, then a vacant position $(a,b)$ is a residue for $R\in\Omega_{U}$ unless it is attacked from above or from its left; see Figure 1 and Remark 3.14 (i). The set of such positions is a well-studied object known as the Rothe diagram of the permutation $\tau$ associated to $R$, and it is a well-known result that the cardinality of the Rothe diagram is equal to the length of $\tau$. In this case, the level function is the length of $\tau$, the quantum Hafnian is the $q$-permanent of a $k\times k$ matrix $A$, and specializing $q$ to q-q gives the $q$-determinant of $A$.
@@
Figure 1. The Rothe diagram of $\underline{3142}\in S_{4}$, with residues marked by stars
Another motivating example (Theorem 3.10) is the case where $U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j<2k\}$ is a set of $\binom{2k}{2}$ roots of type $D_{2k}$ corresponding to the $2$-subsets of the set $2k]:=\{1,2,\ldots,2k\}$. Here, the set $\Omega_{U}$ corresponds to the set of perfect matchings of the set $[2k]$, and a pair $\{a,b\}$ (where $1\leq a<b\leq 2k$) is a residue of $R\in\Omega_{U}$ if and only if we have both $\tau(a)>b$ and $\tau(b)>a$, where $\tau$ is the fixed point free involution corresponding to the perfect matching. The residues in this case are the Rothe diagrams of fixed point free involutions as defined by Hamaker–Marberg–Pawlowski [9, (3.3)], the level function is a well-known statistic on perfect matchings, and the quantum Hafnian is precisely the $q$-Hafnian defined by Jing–Zhang [13]. Specializing $q$ to q-q gives the quantum Pfaffian defined by Strickland [16], and specializing $q$ to +1+1 and to 1-1 gives the Hafnian and Pfaffian of a matrix, respectively. Recall that the Pfaffian, $\text{\rm Pf}(A)$, of a skew-symmetric $2k\times 2k$ matrix $A$ is a polynomial in the entries of $A$ with the property that $(\text{\rm Pf}(A))^{2}=\det(A)$.
@@
In Section 4, we construct many other examples of sets of roots $U$ whose level function gives $\Omega_{U}$ the structure of a quasiparabolic set. Our construction seems to be an unreasonably effective method of producing a quasiparabolic set from a set of roots. For root systems of types $D_{2k}$, $E_{7}$, or $E_{8}$, we can form a quasiparabolic set by taking $U={\Phi_{+}}$ to be the set of all positive roots (Theorem 4.3). We also discuss an abundance of examples of quasiparabolic sets $\Omega_{U}$ where $U=U(\Gamma,p)$ is a set of roots associated to a suitable simple root $\alpha_{p}$ in a simply laced root system. These include examples corresponding to every simple root in every root system of type $A$ or $D$ (Example 4.7, Remark 4.8 (i)), as well as a total of $12$ examples in type $E$ (Example 4.9 (i)). The residue sets for the examples of type $A$ and $D$ have interesting combinatorial interpretations in terms of rook configurations, generalizing the aforementioned motivating examples. One of the examples in type $E_{7}$, associated to the set $U=U(E_{7},2)$, corresponds in a precise way to the $30$ inequivalent $7$-labellings of the Fano plane, and one example in type $E_{8}$, $U=U(E_{8},2)$, corresponds to the $240$ packings of the projective space $PG(3,2)$.
@@
In Section 5, we focus on two important examples arising from the construction of Section 4 in type $E$. The set $U$ in these cases are given by a certain set $U(E_{8},8)$ of $56$ roots in type $E_{8}$ and a certain set $U(E_{7},7)$ of $27$ roots in type $E_{7}$, both of which appear naturally as the set of weights of a suitable minuscule representation. The set $U=U(E_{7},7)$ gives rise to a quasiparabolic set $\Omega_{U}$ whose elements correspond to tritangent planes on a del Pezzo surface of degree 3 and whose Pfaffian gives the well-known invariant cubic polynomial in type $E_{6}$ (Theorem 5.19). The quasiparabolic order on $\Omega_{U}$ endows the tritangent planes with a natural partial order and yields a non-recursive interpretation of the coefficients of the invariant polynomial, both of which are new to our knowledge.
@@
The paper is organized as follows. Section 2 recalls the necessary background regarding the combinatorics of root systems. In Section 3, we introduce the key definitions and summarize the theory of quasiparabolic sets developed by Rains–Vazirani, and we give details of the two motivating examples. Section 4 exhibits many other examples of our construction, including the cases where $U$ is the set of all positive roots of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Section 5 is an in-depth study of the two examples in type $E$ mentioned in the last paragraph. Section 6 discusses some possible directions for further research.
@@ 2.1. Review of root systems
We recall the necessary background concerning root systems in this subsection. While many definitions in this paper work more generally for all (finite) root systems, we will focus on irreducible simply laced root systems, i.e., on the root systems of types $ADE$. We label the vertices in the Dynkin diagrams of root systems as shown in Figure 2 until Section 5.2, where we will switch to a different convention in type $E$ for reasons we will explain.
@@
Figure 2. Dynkin diagrams of irreducible simply laced Weyl groups
Let $\Phi=\Phi(\Gamma)$ be a simply laced root system with Dynkin diagram $\Gamma$. The vertices of $\Gamma$ index the simple roots $\Pi=\{\alpha_{i}:i\in\Gamma\}$, and the rank of $\Phi$ is $\text{\rm rank}(\Phi):=\lvert\Pi\rvert$. The root lattice ${\mathbb{Z}}\Pi$ is the free $\mathbb{Z}$}-module on $\Pi$, and it carries a symmetric $\mathbb{Z}$}-bilinear form $B$ given by
@@
Every root is of the form $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ for some integers $c_{i}$ that are either all all nonnegative or all nonpositive. We say $\alpha$ is positive or negative in these cases, respectively. The sum of the coefficients $\sum_{i\in\Gamma}c_{i}$ is called the height of $\alpha$. Given any $\alpha\in\Phi$, the only multiples of $\alpha$ in $\Phi$ are $\alpha$ and $\alpha$.
@@
A subsystem of a root system $\Phi$ is a subset $\Psi\subseteq\Phi$ that is itself a root system. Equivalently, a subsystem of $\Phi$ is a nonempty subset that is closed under the operations $\{s_{\alpha}:\alpha\in\Psi\}$. The positive roots of $\Phi$ naturally induce a positive system, $\Psi_{+}$, and a simple system, $\Pi_{\Psi}$, in $\Psi$ (in the sense of [10, Section 1.3]; see also [8, Section 2.2]). We have $\Psi_{+}=\Psi\cap\Phi_{+}$, and $\Pi_{\Psi}$ can be characterized as the set the elements of $\Psi_{+}$ that cannot be expressed as nonnegative linear combinations of other elements of $\Psi_{+}$. The elements of $\Pi_{\Psi}$ may or may not be elements of $\Pi$, although it is true that $\Pi\cap\Psi\subseteq\Pi_{\Psi}$. We say that $\Psi$ is of type $\Gamma^{\prime}$ if $\Psi$ is isomorphic to a root system of type $\Gamma^{\prime}$.
The bilinear form $B$ on ${\mathbb{Z}}\Pi$ is nondegenerate. We have $B(\alpha,\beta)\in\{2,1,0,-1,-2\}$ for all roots $\alpha,\beta\in\Phi$ , with $B(\alpha,\beta)=2$ if and only if $\alpha=\beta$ ([10, §6.6]). We say two roots $\alpha$ and $\beta$ are orthogonal if $B(\alpha,\beta)=0.$ The roots orthogonal to a fixed root in any root system $\Phi$ naturally form a subsystem, and it is known that when $\Phi$ has type $E_{8}$, $E_{7}$, $E_{6}$, and $D_{n}$ for $n>4$, such a subsystem has type $E_{7}$, $D_{6}$, $A_{5}$, and $A_{1}+D_{n-2}$, respectively, for all roots in $\Phi$ ([7, Theorem 2.5]). The form $B$ is $W$-invariant in the sense that $B(\alpha,\beta)=B(w(\alpha),w(\beta))$ for all $w\in W$ and all $\alpha,\beta\in\Phi$. In particular, the action of $W$ on $\Phi$ preserves orthogonality.
@@ 2.2. Orthogonal sets of a fixed size
Note that Equation (2.6) indeed defines a $W$-action on $\Phi_{+}^{k}$. Also note that since orthogonal sets in $R$ are linearly independent in the reflection representation $V$, $k$-roots cannot exist for $k>\dim(V)=\text{\rm rank}(\Phi)$. The notion of $k$-roots generalize the notions of $2$-roots and $n$-roots studied in [7, 8], where $n=\text{\rm rank}(\Phi)$; see [7, Remark 3.7] and [8, Section 1].
@@
Let $\Phi$ be an irreducible simply laced root system of rank $n$, and let $W$ be the Weyl group of $\Phi$.
@@
Part (i) is well known, and can be proved similarly to [8, Lemma 3.2] by an inductive argument. The first assertion of (ii) is [8, Lemma 3.2], and the second assertion then follows immediately. ∎

### v2

@@ Abstract.
Let $U$ be a set of positive roots of type $ADE$, and let $\Omega_{U}$ be the set of all maximum-cardinality orthogonal subsets of $U$. For each element $R\in\Omega_{U}$, we define a generalized Rothe diagram whose cardinality we call the level, $\rho(R)$, of $R$. We define the generalized quantum Hafnian of $U$ to be the generating function for $\rho$, regarded as a $q$-polynomial in $U$. In this paper, we study a large number of examples of sets of the form $\Omega_{U}$, and we explore their connections with a variety of widely studied algebraic and combinatorial objects. One of our motivating examples involves a certain set of $k^{2}$ roots in type $D_{2k}$, where the elements of $\Omega_{U}$ can be identified with permutations in $S_{k}$, the generalized Rothe diagrams are the traditional Rothe diagrams associated to permutations, and the generalized quantum Hafnian is the $q$-permanent. More generally, all our examples in types $A$ and $D$ are closely related to perfect matchings and rook configurations, and our examples in type $E$ have applications to del Pezzo surfaces and labellings of Fano planes. Each of our examples also gives rise to a matroid, and a large family of our examples is in a natural correspondence with the equal-rank simply-laced symmetric pairs.
@@
root system, Weyl group, Rothe diagram, quantum Hafnian, symmetric space
@@ 1. Introduction
Let $U$ be a subset of the positive roots of a Weyl group of type $ADE$, and let $\Omega_{U}$ be the set of orthogonal subsets of $U$ of maximum cardinality. For each $R\in\Omega_{U}$, we say that a root $\alpha\in U$ dominates $R$, or that $\alpha$ is $R$-dominating, if $s_{\beta}(\gamma)$ is positive for all $\beta\in R$, where $s_{\beta}$ denotes the reflection associated to $\beta$. We define the generalized Rothe diagram, $\mathscr{D}_{U}(R)$, to be the set of all $R$-dominating roots in $U$, and we call $\rho(R):=|\mathscr{D}_{U}(R)|$ the level of $R$. We define the generalized quantum Hafnian of $U$ to be the element of ${\mathbb{Z}}[U][q]$ given by
@@
Although it is not a priori clear why one might want to study these definitions, which are new to the best of our knowledge, we will show that many choices for $U$ give rise to concise constructions of known algebraic and combinatorial objects, in some cases endowing them with additional structure. For instance, our examples in types $A$ and $D$ are naturally related to perfect matchings and rook configurations (Section 4.2), one family of our examples has a close connection with Cartan’s celebrated classification of Riemannian symmetric spaces and symmetric pairs (Section 4.3), and one of our examples in type $E_{6}$ gives a non-recursive combinatorial interpretation for the 45 terms of a well-known invariant polynomial (Theorem 5.17).
A key motivating example of our constructions is the case where $U=\{\varepsilon_{i}+\varepsilon_{k+j}:1\leq i,j\leq k\}$ is a set of $k^{2}$ roots of type $D_{2k}$ (Theorem 3.14). Here, the set $U$ corresponds to the positions $(i,j)$ on a $k\times k$ grid, and the set $\Omega_{U}$ corresponds to the set of configurations of $k$ nonattacking rooks, or equivalently, the elements of the symmetric group $S_{k}$. If we place the origin at the top left and place a dot at each position corresponding to a root in an element $R\in\Omega_{U}$, then a vacant position $(a,b)$ corresponds to an $R$-dominating root unless it is attacked from above or from its left by a dot; see Figure 1 (where the $R$-dominating roots are marked by stars) and Remark 3.15 (i). The set of such positions is a well-studied object known as the Rothe diagram of the permutation $\tau$ associated to $R$, and it is a well-known result that the cardinality of the Rothe diagram is equal to the length of $\tau$. In this case, the level function is the length of $\tau$, the generalized quantum Hafnian is the $q$-permanent of a $k\times k$ matrix $A$, and specializing $q$ to q-q gives the $q$-determinant of $A$.
@@
Figure 1. The Rothe diagram of $\underline{3142}\in S_{4}$
Another motivating example is the case where $U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j<2k\}$ is a set of $\binom{2k}{2}$ roots of type $D_{2k}$ corresponding to the $2$-subsets of the set $2k]:=\{1,2,\ldots,2k\}$ (Theorem 3.11). Here, the set $\Omega_{U}$ corresponds to the set of perfect matchings of the set $[2k]$ and also to the fixed-point-free involutions of the symmetric group $S_{2k}$. For each $R\in\Omega_{U}$, a pair $\{a,b\}$ (where $1\leq a<b\leq 2k$) corresponds to an $R$-dominating root if and only if we have both $\tau(a)>b$ and $\tau(b)>a$, where $\tau$ is the fixed-point-free involution corresponding to $R$. The generalized Rothe diagrams in this case recover the Rothe diagrams of fixed-point-free involutions as defined by Hamaker–Marberg–Pawlowski [13, (3.3)], the level function is a well-known statistic on perfect matchings, and the generalized quantum Hafnian is precisely the $q$-Hafnian defined by Jing–Zhang [18]. Specializing $q$ to q-q gives the quantum Pfaffian defined by Strickland [21], and specializing $q$ to +1+1 and to 1-1 gives the Hafnian and Pfaffian of a suitable matrix, respectively. Recall that the Pfaffian, $\text{\rm Pf}(A)$, of a skew-symmetric $2k\times 2k$ matrix $A$ is a polynomial in the entries of $A$ with the property that $(\text{\rm Pf}(A))^{2}=\det(A)$.
@@
In Section 4, we construct many other examples of sets of roots $U$ whose level function gives $\Omega_{U}$ the structure of a quasiparabolic set. For root systems of types $D_{2k}$, $E_{7}$, or $E_{8}$, we prove in Section 4.1 that we can construct a quasiparabolic set by taking $U$ to be the set of all positive roots (Theorem 4.3). In Section 4.2, we show that by setting $U$ to be the set $U(\Gamma,p)$ of all positive roots (where $\Gamma$ stands for the Dynkin type of the root system) in which a fixed simple root $\alpha_{p}$ appears with odd coefficient, we obtain an abundance of further examples where $\Omega_{U}$ forms a quasiparabolic set (examples 4.7, 4.9, 4.10, 4.11). In types $A$ and $D$, the generalized Rothe diagrams associated to the elements of $\Omega_{U(\Gamma,p)}$ have interesting combinatorial interpretations in terms of rook configurations, generalizing the aforementioned motivating examples. In type $E$, setting $U=U(E_{7},2)$ and $U=U(E_{8},2)$ yields quasiparabolic sets whose elements are in canonical bijections with labellings of the Fano plane using 7 symbols and 8 symbols, respectively, thus endowing the labellings with natural partial orders induced by the quasiparabolic structure (Remark 4.12).
Besides perfect matchings, rook configurations, and Fano labellings, the quasiparabolic sets $\Omega_{U}$ constructed in sections 4.1 and 4.2 also enjoy several intriguing connections with certain matroids and Riemannian symmetric spaces. In all our examples, every element in $\Omega_{U}$ turns out to be equipped with the structure of a binary matroid that can be used to produce the level generating function of $\Omega_{U}$. For the sets $U=U(\Gamma,p)$, the quasiparabolic sets $\Omega_{U}$ correspond to Riemannian spaces whose associated symmetric pair consists of two simply-laced Lie algebras of the same rank, and $U$ coincides with the set of noncompact roots of the symmetric pair. For example, for $U=U(E_{8},8)$, every element of $\Omega_{U}$ carries the matroid structure of the affine space $AG(2,2)$ (Example 4.13), and $\Omega_{U}$ is a quasiparabolic set corresponding to the symmetric space of type EIX (Example 4.14). The precise correspondences between the sets $\Omega_{U}$, binary matroids, and symmetric spaces are explained in Section 4.3, and we believe that deeper relationships between these objects deserve further study (Remark 6.3).
@@
In Section 5, we focus on two important examples arising from the constructions of Section 4 in type $E$. The sets $U$ in these cases are given by a certain set $U(E_{8},8)$ of $56$ roots in type $E_{8}$ and a certain set $U(E_{7},7)$ of $27$ roots in type $E_{7}$, each of which appears naturally as the set of weights of a suitable minuscule representation. The set $U=U(E_{7},7)$ gives rise to a quasiparabolic set $\Omega_{U}$ whose elements correspond to tritangent planes on a del Pezzo surface of degree 3 and whose generalized Pfaffian gives the well-known invariant cubic polynomial in type $E_{6}$ (Theorem 5.19). The quasiparabolic order on $\Omega_{U}$ endows the tritangent planes with a natural partial order and yields a non-recursive interpretation of the coefficients of the invariant polynomial, both of which are new to our knowledge.
@@
The paper is organized as follows. Section 2 recalls the necessary background regarding the combinatorics of root systems. In Section 3, we introduce the key definitions and summarize the theory of quasiparabolic sets developed by Rains–Vazirani, and we give details of the two motivating examples. Section 4 exhibits many other examples of our construction, as well as their connections to binary matroids and symmetric spaces. Section 5 is an in-depth study of the two examples in type $E$ mentioned in the last paragraph. Section 6 discusses some possible directions for further research.
@@ 2.1. Review of root systems
We recall the necessary background concerning root systems in this subsection. While many definitions in this paper work more generally for all (finite) root systems, we will focus on irreducible simply-laced root systems, i.e., on the root systems of types $ADE$. We label the vertices in the Dynkin diagrams of root systems as shown in Figure 2 until Section 5.2, where we will switch to a different convention in type $E$ for reasons we will explain.
@@
Figure 2. Dynkin diagrams of irreducible simply-laced Weyl groups
Let $\Phi=\Phi(\Gamma)$ be a simply-laced root system with Dynkin diagram $\Gamma$. The vertices of $\Gamma$ index the simple roots $\Pi=\{\alpha_{i}:i\in\Gamma\}$, and the rank of $\Phi$ is $\text{\rm rank}(\Phi):=\lvert\Pi\rvert$. The root lattice ${\mathbb{Z}}\Pi$ is the free $\mathbb{Z}$}-module on $\Pi$, and it carries a symmetric $\mathbb{Z}$}-bilinear form $B$ given by
@@
Every root is of the form $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ for some integers $c_{i}$ that are either all nonnegative or all nonpositive. We say $\alpha$ is positive or negative in these cases, respectively. The sum of the coefficients $\sum_{i\in\Gamma}c_{i}$ is called the height of $\alpha$. Given any $\alpha\in\Phi$, the only multiples of $\alpha$ in $\Phi$ are $\alpha$ and $\alpha$.
@@
A subsystem of a root system $\Phi$ is a subset $\Psi\subseteq\Phi$ that is itself a root system. Equivalently, a subsystem of $\Phi$ is a nonempty subset that is closed under the operations $\{s_{\alpha}:\alpha\in\Psi\}$. The positive roots of $\Phi$ naturally induce a positive system, $\Psi_{+}$, and a simple system, $\Pi_{\Psi}$, in $\Psi$ (in the sense of [15, Section 1.3]; see also [12, Section 2.2]). We have $\Psi_{+}=\Psi\cap\Phi_{+}$, and $\Pi_{\Psi}$ can be characterized as the set of elements of $\Psi_{+}$ that cannot be expressed as nonnegative linear combinations of other elements of $\Psi_{+}$. The elements of $\Pi_{\Psi}$ may or may not be elements of $\Pi$, although it is true that $\Pi\cap\Psi\subseteq\Pi_{\Psi}$. We say that $\Psi$ is of type $\Gamma^{\prime}$ if $\Psi$ is isomorphic to a root system of type $\Gamma^{\prime}$.
The bilinear form $B$ on ${\mathbb{Z}}\Pi$ is nondegenerate. We have $B(\alpha,\beta)\in\{2,1,0,-1,-2\}$ for all roots $\alpha,\beta\in\Phi$ , with $B(\alpha,\beta)=2$ if and only if $\alpha=\beta$ ([15, §6.6]). We say two roots $\alpha$ and $\beta$ are orthogonal if $B(\alpha,\beta)=0.$ The set $\Phi_{\alpha}$ of roots orthogonal to a fixed root $\alpha\in\Phi$ naturally form a subsystem. The form $B$ is $W$-invariant in the sense that $B(\alpha,\beta)=B(w(\alpha),w(\beta))$ for all $w\in W$ and all $\alpha,\beta\in\Phi$. In particular, the action of $W$ on $\Phi$ preserves orthogonality.
Remark 2.1.
It is known that when $\Phi$ has type $E_{8}$, $E_{7}$, $E_{6}$, or $D_{n}$ for $n>4$, for each $\alpha\in\Phi$ the subsystem $\Phi_{\alpha}$ has type $E_{7}$, $D_{6}$, $A_{5}$, or $A_{1}+D_{n-2}$, respectively ([11, Theorem 2.5]).
@@ 2.2. Orthogonal sets of a fixed size
Note that Equation (2.5) indeed defines a $W$-action on $\Phi_{+}^{k}$. Also note that since orthogonal sets in $R$ are linearly independent in the reflection representation $V$, $k$-roots cannot exist for $k>\dim(V)=\text{\rm rank}(\Phi)$. The notion of $k$-roots generalizes the notions of $2$-roots and $n$-roots studied in [11, 12], where $n=\text{\rm rank}(\Phi)$; see [11, Remark 3.7] and [12, Section 1].
@@
Let $\Phi$ be an irreducible simply-laced root system of rank $n$, and let $W$ be the Weyl group of $\Phi$.
@@
Part (i) is well known, and can be proved similarly to [12, Lemma 3.2] by an inductive argument. The first assertion of (ii) is [12, Lemma 3.2], and the second assertion then follows immediately. ∎